A body starts from rest and uniformly accelerated for 10 seconds so that its velocity becomes 100m/s. It then maintains the speed for 10 more seconds. The distance travelled in first 10 seconds to next 10 seconds will be in the ratio
a) 1:1
b) 1:2
c) 1:3
d) 1:4

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

For the first 10 seconds,v=u+at\implies a=\dfrac{100m/s}{10s}=10m/s^2
Hence distance travelled in the time=s_1=\dfrac{1}{2}at^2=500m

For the next 10 seconds, distance travelleds_2=vt=100m/s \times 10s=1000m
Therefores_1:s_2=1:2